OpenStudy (anonymous):

What is the solution to the equation? 1 over 5(2x - 15) + 3 over 5x =4 1 over 4

jhonyy9 (jhonyy9):

1 3 1 - *(2x-15) + -------- = 4 -- 5 5x 4 hence is wrote right ?

OpenStudy (anonymous):

Close. \[\frac{ 1 }{ 5 } \times \left( 2x - 15\right) + \frac{ 3 }{ 5 }x = 4\frac{ 1 }{ 4}\]

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@arabpride